What quantity of 40% acid solution must be mixed with 120 mL of 90% acid solution to produce 70% acid solution?
Answer by richwmiller(17219)

.9*120+.4x=.7(120+x)
9*120+.4x=.7*120+.7x
.2*120=.3x
2*120=3x
2*40=x
80 mL=x